# Statistical Inference for Policy Evaluation with Temporal Difference Learning

### v3

@@ 3 Results
It is known (Polyak and Juditsky, 1992; Fort, 2015) that the TD estimation error of (10) satisfies
[…]
where the asymtoptic covariance matrix $\bm{\Lambda}^{\star}$ is given by
$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}\quad\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (12)
Thus, in the fixed-dimensional setting, the above result implies the rate
@@
which may serve as a benchmark. Our main goal is to demonstrate how the TD algorithm can be used to construct non-asymptotic confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$ in high dimensions. Towards this goal, in the subsequent sections we will establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regions/intervals for $\bm{\theta}^{\star}$ using the estimator $\overline{\bm{\theta}}_{T}$.
@@ 3.1 High-probability bounds
$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}+\tilde{C}T^{-\frac{\alpha+1}{2}}\log\frac{T}{\delta}$
with probability at least $1-\delta$, where $\tilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
The proof can be found in Appendix B.1, where the value of $\tilde{C}$ is given in equation (78). Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs. Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ (in Euclidean distance) at the rate of $O(\sqrt{{\mathsf{Tr}(\bm{\Lambda}^{\star})}/{T}})$ up to a log factor. It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature. To the best of our knowledge, the tightest results were attained by Dalal et al. (2018) for inverse-decaying stepsizes and Li et al. (2024); Samsonov et al. (2024) for constant stepsize as summarized in Table 1. Theorem 3.1 is the first result to achieve tight high-probability convergence guarantees up to log factors using polynomial-decaying stepsizes. Another subtle but consequential difference concerns the restriction on the step size. For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$; specifically, $\eta_{0}$ is required to be sufficiently small such that $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\tilde{C}$, where $\tilde{C}$ is independent of $T$ and $\delta$ but depends on the other parameters of interest; its specific form is shown in (70) and (71) in the Appendix. Though not ideal, this condition is necessary for a Chernoff-type convergence guarantee; we refer to Appendix B.1.3 for a detailed discussion. Nonetheless, our stepsize requirements are considerably weaker than those imposed by Samsonov et al. (2024) and Li et al. (2024), which depend on both $\delta$ and $T$; see Table 1.
@@ 3.2 High-dimensional Berry-Esseen bounds
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))$ $\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$
$+\tilde{C}\;T^{-\frac{1}{2}}+\tilde{C}^{\prime}\;T^{2\alpha-2}$
where $\tilde{C},\tilde{C}^{\prime}$ are problem-related quantities independent of $T$.
On the final rate. By choosing $\alpha=2/3$, we obtain a rate in $T$ of order $O(T^{-1/3})$. In fact, in Theorem B.3 in the Appendix we establish the more refined result that the Berry-Esseen bound (in the convex distance) between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and the Gaussian distribution with (approximately) matching first two moments is of order $\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}$. As $\alpha\to 1$, this rate approaches the optimal rate of $1/\sqrt{T}$. The additional term $\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$ in the above bound, which is ultimately the reason for the final rate $O(T^{-1/3})$, originates from replacing an approximation to the finite sample variance of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ with its asymptotic variance $\bm{\Lambda}^{\star}$.
Comparison with Samsonov et al. (2024). Prior to this work, the best Berry-Esseen bounds for stochastic approximation with polynomial-decaying stepsizes and Polyak-Ruppert averaging are those obtained by Samsonov et al. (2024), thus covering the settings considered here. Specifically, the authors show in their Theorem 3 that when $\eta_{0}\leq\min\{\lambda_{0}(1-\gamma),\frac{1-\gamma}{(1+\gamma)^{2}}\}$ and $\alpha\in[\frac{1}{2},1)$,
@@
Choosing $\alpha=\frac{1}{2}$ leads to a convergence rate of order $O(T^{-\frac{1}{4}})$. In contrast, Theorem 3.2 shows that, by selecting $\alpha=\frac{2}{3}$, we obtain the faster rate $O(T^{-\frac{1}{3}})$. As shown in Appendix B.2.3, this rate is tight for $\alpha\in(\frac{2}{3},1)$. See also Figure 2 below, showing the output of a numerical experiment demonstrating that the choice of $\alpha=\frac{2}{3}$ leads to a smaller approximation error. Though worked out independently, the proofs of our Theorem 3.2 and of Theorem 3 of Samsonov et al. (2024) bear similarities, as they both rely on the general Berry-Esseen bounds for non-linear statistics developed by Shao and Zhang (2022). We obtain sharper rates by a more targeted analysis of the TD settings, while, in comparison, the results of Samsonov et al. (2024) apply more broadly to SA problems. Specifically we (i) adopt a different decomposition of the TD estimation error, (ii) handle in a more refined way certain terms appearing in the bound of Theorem 3.1 of Shao and Zhang (2022) [^2] Specifically, the term $I_{3}$ in Appendix B.2.1. and (iii) carry out a more specialized analysis of the properties of the matrices $\bm{A}_{t}$ and $\bm{A}$ (see, e.g., Lemma A.4 in the Appendix) originating from TD learning, which has allowed us to tighten problem-related coefficient, as well as to work with a substantially larger intial stepsize $\eta_{0}$. This non-trivial sharpening may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions, as, e.g., in Theorem 3.4 of Shao and Zhang (2022).
3.3 Online estimation of the covariance matrix
In the last section, we obtained finite sample bounds on the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and its Gaussian approximation, which features an unknown variance. To construct confidence regions for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$. Two types of approaches have been recently proposed: plug-in estimation (Chen et al., 2020) and bootstraping (White and White, 2010; Hanna et al., 2017; Hao et al., 2021; Ramprasad et al., 2022; Zhu et al., 2023; Samsonov et al., 2024).
Here we propose to use a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$ with superior computational efficiency than existing methodologies and establish its non-asymptotic rate of convergence. We will use
$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t},\quad\text{and}\quad\widehat{\bm{\Gamma}}_{T}=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{T}-\bm{b}_{t})^{\top},$ (14)
as estimators for $\bm{A}$ and $\bm{\Gamma}$ respectively, and set
@@ 3.3 Online estimation of the covariance matrix
Crucially, because of the linear nature of the problem, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online manner. See Algorithm 1 in Appendix D.1 for a detailed description of the procedure, which has time and space complexities of order $O(Td^{4})$ and $O(d^{4})$, respectively. The estimator $\widehat{\bm{\Lambda}}_{T}$ also enjoys good statistical properties. This is made precise by the following Gaussian comparison result, whose proof can be found in Appendix B.3, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem B.4 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation (and without affecting the rate in $T$).
@@
Above, the term $o(T^{-\frac{1}{3}})$ depends on all the parameters of the problem (and, in particular, to $\alpha$) but, as a function of $T$ alone, is of smaller order than $T^{-\frac{1}{3}}$. See Appendix B.3 for the proof and details.
Comparison with Chen et al. (2020). In different settings, Chen et al. (2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function. In their work, they consider the slightly different plug-in estimator $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}_{T}^{\top}$, where $\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}$. The authors proved that the estimation error of $\widehat{\bm{\Lambda}}_{T}^{\prime}$ is bounded in expectation as
@@
It is noteworthy that the online Algorithm 1 takes advantage of the linearity of the problem; in general, potentially non-linear settings, replacing $\bm{\theta}_{t}$ in $\widehat{\bm{\Gamma}}_{T}^{\prime}$ by $\overline{\bm{\theta}}_{T}$ in $\widehat{\bm{\Gamma}}_{T}$ would require storing all the samples throughout the iterative trajectory, resulting in a larger space complexity.
Comparison with Samsonov et al. (2024).
Leveraging the multiplier-bootstrap technique proposed by Zhu et al. (2023), Samsonov et al. (2024) constructs a randomized estimator $\overline{\bm{\theta}}_{T}^{(b)}$ based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$. Their Theorems 4 and 5 imply that, in our notation,
@@
Compared with this bootstrap-based procedure, our methodology is advantageous in the following ways: (i) the upper bound in Theorem 3.3 vanishes, as a function of $T$, at the rate of $O(T^{-\frac{1}{3}})$, which is faster than the $O(T^{-\frac{1}{4}}\log T)$ rate from the bound (18). (ii) The leading term in the upper bound in Theorem 3.3 entails a smaller problem-related coefficient than that of (18). (iii) Perhaps most importantly, in order to estimate the bootstrap distribution $\sqrt{T}(\overline{\bm{\theta}}_{T}^{(b)}-\bm{\theta}^{\star})|\mathcal{S}_{T}$ with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories. Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$. When $d\ll T$, this amounts to a much higher computational burden than our online Algorithm 1.

### v4

@@ 3 RESULTS
It is known (Polyak and Juditsky,, 1992; Fort,, 2015) that, in fixed-dimensional settings, the TD estimation error of (11) satisfies
$\sqrt{T}\,\overline{\bm{\Delta}}_{T}\xrightarrow{d}\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}),$ (12)
where the asymptotic covariance matrix $\bm{\Lambda}^{\star}$ is given by
$\bm{\Lambda}^{\star}=\bm{A}^{-1}\bm{\Gamma}\bm{A}^{-\top},\quad\text{with}$ (13)
$\bm{\Gamma}=\mathbb{E}[(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})(\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t})^{\top}].$ (14)
Thus, the above result implies the familiar paranetric rate
@@
which may serve as a benchmark. Below, we will extend these guarantees to a high-dimensional scenario by demonstrating an explicit dependence on all model-related parameters including, in particular, the dimension of the feature space. Our main goal is to show how the TD algorithm can be used to construct non-asymptotic confidence regions and simultaneous confidence intervals for $\bm{\theta}^{*}$ in high dimensions. To that effect, we will establish a high-probability convergence guarantee for $\overline{\bm{\Delta}}_{T}$, develop high-dimensional Berry-Esseen bounds over the class of convex sets, and finally build confidence regionsfor $\bm{\theta}^{\star}$ using $\overline{\bm{\theta}}_{T}$ and a novel, online estimator of the asymptotic variance.
@@ 3.1 High-probability bounds
$\|\overline{\bm{\Delta}}_{T}\|_{2}$ $\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{\Lambda}^{\star})}{T}\log\frac{1}{\delta}}+\widetilde{C}(T^{-\frac{\alpha+1}{2}}+T^{\frac{\alpha}{2}-1})\log\frac{T}{\delta}$
with probability at least $1-\delta$, where $\widetilde{C}$ is a problem-related quantity independent of $T$ and $\delta$.
The proof of this result can be found in Appendix B.1, where the value of $\widetilde{C}$ is given in equation (81). Throughout the paper, for the sake of readability, we have opted to express unwieldy problem-dependent quantities that do not directly depend on $T$ in implicit form; their values can be tracked in the proofs. Accordingly, when writing our finite-sample bounds, we will first present the leading terms in $T$.
In a nutshell, Theorem 3.1 states that $\overline{\bm{\theta}}_{T}$ converges to $\bm{\theta}^{*}$ (in Euclidean distance) at the rate of order $O(\sqrt{{\mathsf{Tr}(\bm{\Lambda}^{\star})}/{T}})$, up to a log factor. It is worth comparing the high-probability sample complexity bound of Theorem 3.1 with those in the literature. To the best of our knowledge, the tightest results were attained by Dalal et al., (2018) for inverse-decaying stepsizes and Li et al., 2024b ; Samsonov et al., 2024b for constant stepsize as summarized in Table 1. Theorem 3.1 is the first result to achieve tight high-probability convergence guarantees up to log factors using polynomial-decaying stepsizes. Another subtle but consequential difference concerns the restriction on the step size. For our bound to hold, the choice of the initial stepsize $\eta_{0}$ depends on $\delta$ – specifically, $\eta_{0}$ is required to be sufficiently small such that $\eta_{0}\sqrt{\log{1}/(\delta\eta_{0}})\leq\widetilde{C}$, where $\widetilde{C}$ is independent of $T$ and $\delta$ but depends on the other parameters of interest (its specific form is shown in (73) and (74) in the Appendix). Though not ideal, this condition is necessary for a Chernoff-type convergence guarantee; we refer to Appendix B.1.3 for a detailed discussion. Nonetheless, our stepsize requirements are considerably weaker than those imposed by Samsonov et al., 2024b and Li et al., 2024b , which depend on both $\delta$ and $T$; see Table 1.
@@ 3.2 High-dimensional Berry-Esseen bounds
$d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}+\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}+\widetilde{C}\;T^{-\frac{1}{2}}+\widetilde{C}^{\prime}\;T^{2\alpha-2}$
where $\widetilde{C},\widetilde{C}^{\prime}$ are problem-related quantities independent of $T$.
On the final rate. By choosing $\alpha=2/3$, we obtain a rate in $T$ of order $O(T^{-1/3})$. In fact, in Theorem B.3 in the Appendix we establish the more refined result that the Berry-Esseen bound (in the convex distance) between the distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and a Gaussian distribution with (approximately) matching first two moments is of order $\sqrt{\frac{\eta_{0}}{(1-\alpha)(1-\gamma)\lambda_{0}}}\mathsf{Tr}(\bm{\Gamma})\|\bm{\Gamma}^{-1}\|T^{-\frac{\alpha}{2}}$. As $\alpha\to 1$, this rate approaches the optimal rate of $1/\sqrt{T}$. The additional term $\frac{\sqrt{d\mathsf{cond}(\bm{\Gamma})}}{(1-\gamma)\lambda_{0}\eta_{0}}T^{\alpha-1}$ in the above bound, which is ultimately the reason for the final rate $O(T^{-1/3})$, originates from replacing an approximation to the finite sample variance of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ with its asymptotic variance $\bm{\Lambda}^{\star}$.
Comparison with Samsonov et al., 2024a . Prior to this work, the best Berry-Esseen bounds for stochastic approximation with polynomial-decaying stepsizes and Polyak-Ruppert averaging are those obtained by Samsonov et al., 2024a , thus covering the settings considered here. Specifically, the authors show in their Theorem 3 that when $\eta_{0}\leq\min\{\lambda_{0}(1-\gamma),\frac{1-\gamma}{(1+\gamma)^{2}}\}$ and $\alpha\in[\frac{1}{2},1)$,
@@
Choosing $\alpha=\frac{1}{2}$ leads to a convergence rate in $T$ of order $O(T^{-\frac{1}{4}})$. In contrast, Theorem 3.2 shows that, by selecting $\alpha=\frac{2}{3}$, we obtain the faster rate $O(T^{-\frac{1}{3}})$. As demonstrated in Appendix B.2.3, this rate is tight for $\alpha\in(\frac{2}{3},1)$, since the $O(T^{\alpha-1})$ term is inevitable. See also Figure 2 below, showing the output of a numerical experiment indicating that, in fact, the choice of $\alpha=\frac{2}{3}$ leads to a smaller approximation error. Though worked out independently, the proofs of our Theorem 3.2 and of Theorem 3 of Samsonov et al., 2024a bear similarities, as they both rely on the general Berry-Esseen bounds for non-linear statistics developed by Shao and Zhang, (2022). We obtain sharper rates by a more targeted analysis of the TD settings, while, in comparison, the results of Samsonov et al., 2024a apply more broadly to SA problems. Specifically we (i) adopt a different decomposition of the TD estimation error, (ii) handle in a more refined way certain terms appearing in the bound of Theorem 3.1 of Shao and Zhang, (2022) [^2] Specifically, the term $I_{3}$ in Appendix B.2.1. and (iii) carry out a more specialized analysis of the properties of the matrices $\bm{A}_{t}$ and $\bm{A}$ (see, e.g., Lemma A.4 in the Appendix) originating from TD learning, which has allowed us to tighten problem-related coefficient, as well as to work with a substantially larger initial stepsize $\eta_{0}$. This non-trivial sharpening may be potentially generalized to analyze stochastic gradient algorithms applied to non-linear functions, as, e.g., in Theorem 3.4 of Shao and Zhang, (2022).
3.3 Online estimation of the covariance
In the last section, we obtained finite sample bounds on the convex distance between $\sqrt{T}\overline{\bm{\Delta}}_{T}$ and its Gaussian approximation, which features an unknown variance. To construct confidence regions for $\bm{\theta}^{\star}$, it is necessary to estimate the covariance matrix $\bm{\Lambda}^{\star}$. Two types of approaches have been recently proposed: plug-in estimation (Chen et al.,, 2020) and bootstrapping (White and White,, 2010; Hanna et al.,, 2017; Hao et al.,, 2021; Ramprasad et al.,, 2022; Zhu et al.,, 2023; Samsonov et al., 2024a, ).
Here we propose to use a computationally feasible online plug-in estimator for $\bm{\Lambda}^{\star}$ with superior computational efficiency than existing methodologies and establish its non-asymptotic rate of convergence. Specifically, we consider the following
$\overline{\bm{A}}_{T}=\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t},\quad\text{and}$ (16)
$\widehat{\bm{\Gamma}}_{T}=\sum_{t=1}^{T}w_{t,T}(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}\overline{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}$ (17)
as estimators for $\bm{A}$ and $\bm{\Gamma}$ respectively, where the weights $\{w_{t,T}\}_{1\leq t\leq T}$ in $\widehat{\bm{\Gamma}}_{T}$ are chosen as
$w_{t,T}:=\frac{t}{\sum_{t=1}^{T}t}=\frac{2t}{T(T+1)},$
following the intuition that as $t$ grows larger, $\overline{\bm{\theta}}_{t}$ is a more accurate estimator of $\bm{\theta}^{\star}$ and hence the corresponding term should be assigned higher weights. We then consider
@@ 3.3 Online estimation of the covariance
Crucially, the estimator $\widehat{\bm{\Lambda}}_{T}$ can be evaluated in an online fashion. See Algorithm 1 in Appendix D.1 for a detailed description of the procedure, which does not require a priori knowledge of the total round of iterations $T$, and has time and space complexities of order $O(Td^{2}+d^{3})$ and $O(d^{2})$, respectively. The estimator $\widehat{\bm{\Lambda}}_{T}$ also enjoys nice statistical properties. This is made precise by the following Gaussian comparison result, showing that the unknown asymptotic variance $\bm{\Lambda}^{\star}$ in the Berry-Esseen bound of Theorem B.4 can be replaced by $\widehat{\bm{\Lambda}}_{T}$ without compromising the accuracy of the approximation (i.e., without affecting the rate in $T$). The proof of this result can be found in Appendix B.3.
@@
Above, the term $o(T^{-\frac{1}{3}})$ depends on all the parameters of the problem (and, in particular, on $\alpha$) but, as a function of $T$ alone, it is of order $O(T^{-\frac{1}{3}})$. See Appendix B.3 for the proof and details.
Comparison with Chen et al., (2020). In different settings, Chen et al., (2020) considered the problem of carrying out statistical inference for the true model parameters using stochastic gradient descent and assuming a strongly convex and smooth population loss function. In their work, they consider the slightly different plug-in estimator $\widehat{\bm{\Lambda}}_{T}^{\prime}=\overline{\bm{A}}_{T}^{-1}\widehat{\bm{\Gamma}}_{T}^{\prime}\overline{\bm{A}}_{T}^{\top}$, where $\widehat{\bm{\Gamma}}_{T}^{\prime}:=\frac{1}{T}\sum_{t=1}^{T}(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})(\bm{A}_{t}{\bm{\theta}}_{t}-\bm{b}_{t})^{\top}$. The authors proved that the estimation error of $\widehat{\bm{\Lambda}}_{T}^{\prime}$ is bounded in expectation as
@@
Comparison with Samsonov et al., 2024a .
Leveraging the multiplier-bootstrap technique proposed by Zhu et al., (2023), Samsonov et al., 2024a construct a randomized estimator $\overline{\bm{\theta}}_{T}^{(b)}$ based on the samples $\mathcal{S}_{T}:=\{s_{t},s_{t}^{\prime}\}_{t=1}^{T}$. Their Theorems 4 and 5 yield that
@@
Compared with this bootstrap-based procedure, our methodology is advantageous in the following ways: (i) the upper bound in Theorem 3.3 vanishes, as a function of $T$, at the rate of $O(T^{-\frac{1}{3}})$, which is faster than the $O(T^{-\frac{1}{4}}\log T)$ rate from the bound (21). (ii) The leading term in the upper bound in Theorem 3.3 entails a smaller problem-related coefficient than that of (21). (iii) Perhaps most importantly, in order to estimate the bootstrap distribution with an $O(T^{-\frac{1}{4}})$ error rate, it is necessary to store the sample trajectory $\mathcal{S}_{T}$ and generate $O(T^{\frac{1}{2}})$ perturbed trajectories. Consequently, the bootstrap procedure has a time complexity of order $O(T^{\frac{3}{2}}d^{2})$ and a space complexity of order $O(T)$. This amounts to a much higher computational burden than our online Algorithm 1.